how to breakdown the cost of a dollar cup of coffee into percentages representing a pie chart where percentages repesent how much of each dollar goes for rent.paper,food,utilities,payroll,advertising,profit,etc. Answer by stanbon(75887) 
 Example:
cup of coffee: 1.50
orange juce: 2.00
danish: 2.50
-------------------------
Add them up to get 6.00
--------------------------
Convert each to a percentage:
coffee: 1.50/6.00 = 25%
orange juice: 2.00/6.00 = 33 1/3%
danish: 2.50/6.00 = 41 2/3%
--------------------------------
Draw the pie and estimate the percentage sectors; then label them.
=======================================
Cheers,
Stan H.